# Is Your LLM Outdated? Evaluating LLMs at Temporal Generalization

### v2

@@ Abstract
The rapid advancement of Large Language Models (LLMs) highlights the urgent need for evolving evaluation methodologies that keep pace with improvements in language comprehension and information processing. However, traditional benchmarks, which are often static, fail to capture the continually changing information landscape, leading to a disparity between the perceived and actual effectiveness of LLMs in ever-changing real-world scenarios. Our study examines temporal generalization, which includes the ability to understand, predict, and generate text relevant to past, present, and future contexts, revealing significant temporal biases in LLMs. We propose an evaluation framework, for dynamically generating benchmarks from recent real-world predictions. Experiments demonstrate that LLMs struggle with temporal generalization, showing performance decline over time. These findings highlight the necessity for improved training and updating processes to enhance adaptability and reduce biases. Our code, dataset and benchmark are available at https://github.com/FreedomIntelligence/FreshBench.
@@ 1 Introduction
Proprietary large language models (LLMs) such as GPT-4 [1], Bard [2], Claude [3], and PaLM [4] has been paralleled by efforts in the open-source community to democratize LLMs. Meanwhile, the rapid push for LLM democratization [5] emphasizes the necessity for dynamic and robust benchmarks that accurately reflect their evolving capabilities. Traditional benchmarks, broadly classified into knowledge-based assessments and open-dialogue evaluations, each present unique challenges. Knowledge-based assessments, exemplified by MMLU [6] and C-Eval [7], are prone to data manipulation, raising concerns about their real-world applicability and the risk of data leakage. Open-dialogue evaluations like MT-Bench [8] and Alpaca-Eval [9], rely on subjective human or model-based judgments, making their outcomes susceptible to the influence of dialogue structure, and thereby, potentially compromising assessment validity.
Given the limitations of traditional methods, researchers are exploring new ways to measure model performance more comprehensively. Recently, several methods have been proposed to evaluate models at different times and in various contexts[10, 11, 12]. TimeQA[13] evaluates a model’s ability to handle temporal evolution, while Situated QA[14] focuses on a model’s performance in specific contexts. StreamingQA[15] introduces the concept of continuous information flow, while RealTimeQA[11] and FreshQA[12] emphasize real-time and up-to-date information processing capabilities. However, existing methods may be prone to data leakage during their evaluation processes and typically focus on current or past data, lacking exploration of future event prediction capabilities.
@@
We believe that the evaluation of LLMs should be conducted in scenarios that are resistant to hacking manipulation, with results that are objective and free from evaluator biases [16]. Additionally, time as a dynamic dimension, can also ensure robust and relevant evaluations. This paper explores an alternative way to evaluate large language models, i.e., on temporal generalization where time is considered as an important aspect. By doing so, the inaccessibility of future content would largely reduce the data contamination issue. Moreover, a retrospective perspective on time could be eneficial for temporal bias analysis for LLMs.
@@
To achieve this objective, we explore two specific scenarios designed to measure model capabilities without data leakage and to assess model performance on tasks relevant to both past and future contexts. The first scenario involved generating text from recent sources, such as arxiv papers, news articles, or new books, and Wikipedia. We expected the models to generalize to these diverse, newly generated texts with high language likelihood, thereby supporting the concept of "intelligence as compression" in a manner free from data leakage. The second scenario focused on predicting future events, suggesting that a superior model should be able to understand the present context, incorporate world knowledge, and generalize to future scenarios. In summary, the evaluation at temporal generalization offers two main benefits: it eliminates the possibility of data leakage through temporal boundaries, and its results can be objectively assessed without reliance on subjective interpretations.
@@
Our contributions are as follows: 1) We thoroughly define and quantify temporal generalization and bias, establishing a basis for understanding and assessing the capabilities of LLMs over time. 2) We proposed an benchmark Freshbench to test the capabilities of LLMs in predicting future events. This benchmark is dynamically updated to reflect the latest data, ensuring that our evaluations remain relevant and accurate in the face of rapidly evolving data environments. It experimentally assesses performance of existing LLMs across various timeframes, providing insights into their historical and ongoing performance stability and adaptability.
@@ 2 Conceptualization of Temporal Generalization
Figure 1: Schematic figure of temporal generalization.
In this section, we will conceptualize Temporal Generalization in Sec. 2.1 and further measure it in 2.2. From a retrospective perspective in the Temporal Generalization, we also conceptualize Temporal Biases in Sec. 2.3. In general, prior to the cutoff time, a LLM might exhibit a Nostalgia bias that is skewed toward past dates (i.e., aligning better with historical data) or a Neophilia bias that is skewed toward future dates ( favoring more recent information). However, after the cutoff time, the model’s understanding of future trends declines, leading to temporal degeneration.
@@ 2.1 Temporal Generalization
LLMs face the ongoing challenge of remaining applicability as new text evolves. Consequently, it is crucial for these models to not only understand and generate text based on seen (e.g., past) data but also to anticipate and adapt to future data. This inspires temporal generalization as below.
@@
Temporal Generalization refers to the LLMs’ ability to accurately understand, predict, and generate text relevant to past, present, and future contexts.
@@ 2.2 Assessing Temporal Generalization
Based on definition in Sec. 2.1, we turn to assess temporal generalization through the lens of 1) I: compression intelligence in Sec. 2.2.1 or II): prognostication prediction in Sec. 2.2.2.
2.2.1 Case I: Compression Intelligence over Time
@@
$\small BPC(T)=\frac{-\sum_{i=1}^{N}\log p(w_{i}|w_{1},\ldots,w_{i-1})}{\text{len-utf-8}(T)}$ (1)
@@
2.2.2 Case II: Prognostication Prediction
LLMs could performs prognostication prediction, as seen in [18] which employed GPT-3 in a few-shot setting to tackle binary questions from the Metaculus platform. Further, [19] enriched the discourse by introducing a comprehensive dataset that includes forecasting questions alongside a news corpus. This approach enhances model training by incorporating varied and relevant data contexts. Moreover, [20] showcased how integrating news content can significantly improve the accuracy of predictions in real-time scenarios.
Building on this foundation, our research seeks to explore the temporal generalization capabilities of LLMs: how well they adapt and remain accurate as new data emerges and contexts evolve?. As an example of using LLMs to predict political outcomes, we gather new factual data about future events, such as "What will President Biden’s approval rating be as of 7 June 2024, according to FiveThirtyEight?". To ensure objectivity in our predictions, answers are typically formatted as ranges within a prediction market scenario. For instance, President Biden’s approval rating might be categorized into intervals such as lower than 34%, at least 34.0%, but less than 36.0%, at least 36.0%, but less than 38.0%, up to 46.0% or higher.
@@ 2.2.2 Case II: Prognostication Prediction
$\small Acc=\frac{N_{\text{correct}}}{N_{\text{total}}}$ (2)
@@
2.3 Temporal Bias in a Retrospective View
Temporal biases are multifaceted in previous literature. For instance, temporal bias can refer to the assumption of an incorrect sequence of events that misleads our reasoning about causality [21] or user preferences over time in point-of-interest recommendations [22]. Recently, biases have raised more challenges in pre-trained language models [23], particularly focusing on under-represented societies. [24] examines lookahead bias in pre-trained language models, providing insights into how these biases affect model performance. In the context of LLMs, temporal bias refers to the preferences of LLMs over time.
There exists two biases, namely Nostalgia Bias (see Definition 2.2) and Neophilia Bias (see Definition 2.3), which reflect two contrasting tendencies in how these models might predict or generate text about the past, present, or future. A visual representation of these biases is provided in Fig. 1, where the trends associated with Nostalgia and Neophilia biases are plotted over time.
Definition 2.2.
Definition 2.3.
Challenges in Temporal Generalization Predicting future text with LLMs while navigating Nostalgia Bias and Neophilia Bias presents distinct challenges. Nostalgia Bias can lead to outdated predictions, while overcorrection towards Neophilia Bias might result in speculative predictions disconnected from past trajectories. Balancing these biases is crucial to leverage historical and contemporary data effectively, ensuring future predictions are informed by past patterns and open to new possibilities. This requires careful adjustment of the model’s training data and integration of up-to-date information.
@@ 3 Temporal Generalization through the Lens of Compression Intelligence
Temporal generalization in language models evaluates their ability to maintain performance over time, even as data evolves. This section discusses how various models handle data from different periods, showcasing their performance stability or volatility.
@@ 3.1.1 Protocol for Quantitative Measurement
Our system aggregates data from various platforms, including Yahoo for financial news, BBC for political insights, Reddit for discussions, Wattpad for literature, Wikipedia for encyclopedic updates, arXiv for academic research, GitHub for software trends, and Quora for diverse questions and answers. Utilizing a Python-based crawling framework, detailed in our public repository, it’s designed for adaptability and efficient data collection, with Playwright addressing navigation and specific site challenges. We also did some basic data pre-processing, due to space limitation, details in Appendix B. An overview of the data volume and the average entry length from a single crawl is provided in Appendix C.
Data sources of Fresh Corpora We selected a set of models due to their diversity and significance in advancing the capabilities of NLP systems. They are not only benchmarks of their respective sizes, ranging from 3 billion to 13 billion parameters, but also exemplify the rapid progression in NLP technology. The models originate from diverse sources, showcasing the collaborative and competitive spirit driving innovation in the industry.
Experimental Settings To further delve into the model’s performance dynamics over varying text lengths, we truncate the input text to the first 5000 characters and compute the sum of loss. The loss values are then accumulated to derive their sum, providing insight into the model’s efficiency at different stages of text comprehension. Throughout this experiment, approximately 300 GPU hours on NVIDIA A800 were consumed.
@@ 3.1.2 Observation of Temporal Bias in Compression Intelligence
To comprehensively explore the temporal generalization capabilities of models, we plot the BPC curve of fresh text over time. In text data before release, a negative slope indicates the presence of Neophilia Bias. Conversely, a positive slope would suggest that the model’s performance is decreasing, which shows Nostalgia Bias.
We sampled arXiv papers from 2015 to 2024 and applied smoothing techniques to the data because the number of samples was not sufficient to completely stabilize the average language likelihood. Nonetheless, a slight downward trend in model BPC can be observed, which is noteworthy. What is more, models with initially higher BPC showed a larger percentage decrease in BPC (as seen in models like OPT-2.7B and PHI-2 in Figure  3). For models with lower initial BPC, the decrease trend is less pronounced. Taking arXiv text as an example, BPC decreases over time, LLMs exhibit a Neophilia bias rather than a Nostalgia bias, capturing the present more effectively than the past.
Turning to BBC news text, the observed trends in BPC from 2020 to 2023, depicted in Fig. 3, provide insight into how well these models process and understand news content over a span of several years. We note a pronounced increase in BPC across the observed period, which suggests a diminishing adaptability of the models to newer data, indicative of a Nostalgia bias. This trend may stem from the dynamic and evolving nature of news narratives. Detailed result for various sources are in E
Figure 2: BPC on arXiv text from 2015.9 to 2024.1 (smoothed)
Figure 3: BPC on BBC news text from 2020 to 2023, with three-month intervals (smoothed)
3.2 Temporal Generalization
@@ 3.2.1 Protocol for Quantitative Measurement
Temporal Bias Index (TBI) For each model in a certain period, we record $BPC_{i}$ at $n$ time points $t_{i}$, forming $n$ data points $(t_{i},p_{i})$, where $i=1,\ldots,n$. We then fit a linear model:
$\small BPC_{i}=a*t_{i}+b$ (3)
where Temporal Bias Index (TBI) is defined as the coefficient $a$. TBI is crucial for understanding the direction and rate of change in model performance over time.
To quantify model temporal generalization, we computed the BPC for various models on the Wiki and BBC datasets and evaluated their performance using this index. Specifically, We established a base BPC using data from the six months prior to each model’s release. After that, we measured the changes in BPC at 3, 6, 9 and 12 months post-release relative to the base BPC; and assessed the overall trend in model performance using TBI. The specific results are detailed in Appendix D.
Table 1: Correlation between Model Attributes and Their Performance Decay Across Various Tasks on Wikipedia and BBC Benchmarks
@@ 3.2.2 Observation
Models perform stably on Wikipedia but show more variation on BBC. In the Wikipedia dataset, models exhibit stable performance, demonstrating their ability to handle well-structured and formal text. However, larger models in the Wikipedia dataset deteriorate faster (correlation coefficient of 0.504), likely due to overfitting. In contrast, the BBC dataset shows significant variations in model performance, indicating insufficient adaptability to dynamic and informal text. From 2020 to 2024, the overall BPC for BBC texts shows an increasing trend, indicating gradual performance deterioration. Individual models show considerable differences in performance on new data; for instance, the Yi, LLaMA2, and Baichuan2 series adapt relatively well to new texts, while Phi-1.5 and OPT series exhibit poorer generalization (see Appendix E). These findings indicate that in dynamic environments, different models have significantly varying adaptation capabilities.
Positive correlation between model attributes and performance decay. (see Tab. 1). There is a positive correlation between model attributes (such as model size and release date) and performance decay. A positive BPC slope indicates an increase in BPC over time, reflecting performance deterioration. Larger model sizes may lead to rapid performance decay, suggesting a trade-off between capability and flexibility. In dynamic environments (such as the BBC dataset), the impact of model size on performance decay is smaller. Additionally, newer models show higher correlations, meaning that despite their advanced capabilities, they may deteriorate faster. These results emphasize the importance of designing models that can maintain robustness over time and across various data environments. As machine learning applications grow in complexity and scale, temporal generalization will become increasingly critical.
4 Temporal Generalization through the Lens of Prognostication Prediction
Advantages of Prognostication Prediction over Compression Intelligence Compression Intelligence is limited by their inapplicability to closed-source models, where key indicators like logits or losses are inaccessible. This exclusion narrows comparative analysis and hinders a comprehensive understanding of the language technology landscape. Additionally, language likelihood can be biased by models finely tuned to specific formats or domains, skewing overall evaluation and presenting a distorted view of a model’s general performance.
This motivated us to research temporal generalization and bias from the perspective of future prediction. In this section, we investigate tempera generalization (in Sec. 4.1) and biases (Sec. 4.2) through the Lens of Prognostication Prediction.

### v3

# Is Your LLM Outdated? A Deep Look at Temporal Generalization

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has led to the development of benchmarks that consider temporal dynamics, however, there remains a gap in understanding how well these models can generalize across temporal contexts due to the inherent dynamic nature of language and information. This paper introduces the concept of temporal generalization in LLMs, including bias in past and future generalizations. Then we introduce FreshBench, a new evaluation framework that employs fresh text and event prediction for assessing LLMs’ temporal adaptability, ensuring the evaluation process free from data leakage and subjective bias. The experiment shows significant temporal biases and a decline in performance over time. Our findings reveal that powerful models, while initially superior, tend to decline more rapidly in future generalization. Additionally, powerful open-source models demonstrate better long-term adaptability compared to their closed-source counterparts. Our code is available at https://github.com/FreedomIntelligence/FreshBench
@@ 1 Introduction
Proprietary large language models (LLMs)  OpenAI (2023); GoogleAI (2023); Anthropic (2023); Anil et al. (2023) has been paralleled by efforts in the open-source community to democratize LLMs. Meanwhile, the rapid advancement of LLM democratization Touvron et al. (2023) emphasizes the necessity for dynamic and robust benchmarks that accurately reflect their evolving capabilities. Traditional benchmarks, broadly classified into knowledge-based assessments and open-dialogue evaluations, each present unique challenges. Knowledge-based assessments, exemplified by MMLU Hendrycks et al. (2021) and C-Eval Huang et al. (2023), are prone to data manipulation, raising concerns about their real-world applicability and the risk of data leakage Wei et al. (2023); Golchin and Surdeanu (2024). Open-dialogue evaluations like MT-Bench Zheng et al. (2023) and Alpaca-Eval Li et al. (2023a), rely on subjective human or model-based judgments, making their outcomes susceptible to the influence of dialogue structure, and thereby, potentially compromising assessment validity.
Given the limitations of traditional methods, researchers are increasingly developing new approaches to evaluate models in diverse and dynamic contexts. For instance, recent efforts have introduced methods that assess models’ abilities to handle temporal evolution, contextual understanding, and real-time information updates Fatemi et al. (2024); Kasai et al. (2024); Vu et al. (2023); Chen et al. (2021); Zhang and Choi (2021); Liška et al. (2022). While methods like RealTimeQA and FreshQA focus on continuously updating knowledge for real-time processing Kasai et al. (2024); Vu et al. (2023), others like LatestEval Li et al. (2023b) emphasize recent text-based comprehension. In contrast, our work shifts the focus to evaluating models’ world understanding through prediction, using a fact-based approach that minimizes bias and ensures objectivity in assessment. This method enhances robustness against data contamination and focuses on the model’s capacity for accurate, real-world event prediction.
@@
We believe that the evaluation of LLMs should be conducted in scenarios that are resistant to hacking manipulation, with results that are objective and free from evaluator biases Chen et al. (2024). Additionally, time as a dynamic dimension, can also ensure robust and relevant evaluations. This paper explores an alternative way to evaluate large language models, i.e., on temporal generalization where time is considered as an important aspect. This approach significantly reduces the issue of data contamination by limiting access to future content. Moreover, a retrospective perspective on time could be beneficial for temporal bias analysis for LLMs.
@@
To achieve this goal, we design two scenarios. The first is to have the model generate text based on recent data such as arXiv papers, news articles, new books, and Wikipedia content. An excellent model should be able to generalize these new texts with high language likelihood without data leakage. The second scenario focuses on future event prediction. An excellent model should be able to understand the current context, integrate world knowledge, and effectively generalize to future scenarios. In these two scenarios, temporal generalization evaluation has two main advantages: eliminating the possibility of data leakage through temporal boundaries and enabling objective evaluation results without relying on subjective interpretations.
@@
Our contributions are as follows: 1) We define and quantify temporal generalization and bias, establishing a basis for understanding and assessing the capabilities of LLMs over time. 2) We proposed a benchmark Freshbench to test the capabilities of LLMs in predicting future events. This benchmark is dynamically updated to reflect the latest data, ensuring that our evaluations remain relevant and accurate in the face of rapidly evolving data environments. It experimentally assesses the performance of existing LLMs across various timeframes, providing insights into their historical and ongoing performance stability and adaptability.
@@ 2 Conceptualization of Temporal Generalization
In this section, we will first conceptualize Temporal Generalization, then Temporal Biases and Temporal Degeneration in Sec. 2.1. We will discuss ways to measure /textitTemporal Biases and Degeneration in 2.2.
@@ 2.1 Temporal Generalization
LLMs face the ongoing challenge of remaining applicable as new text evolves. Consequently, it is crucial for these models to not only understand and generate text based on seen (e.g., past) data but also to anticipate and adapt to future data. This inspires temporal generalization as below.
Figure 1: Schematic figure of temporal generalization. In general, prior to the cutoff time, an LLM might exhibit a Nostalgia bias that is skewed toward past dates (i.e., aligning better with historical data) or a Neophilia bias that is skewed toward future dates ( favoring more recent information). However, after the cutoff time, the model’s understanding of future trends declines, leading to temporal degeneration.
@@
Temporal Generalization refers to the LLMs’ ability to align with the contexts of the past, present, and future.
@@
2.1.1 For Models: Define Past, Present and Future
Figure 2: Schematic Diagram of 3 Periods of Models
To distinguish the past and the future, we define the cutoff time as the release date of the model. For instance, if an LLM is released in January 2024, then all events or information before January 2024 are considered part of the past, while everything after January 2024 is considered the future. Although the data after the cutoff time is not strictly future data, this approach ensures no data leakage for the model. As is shown in Figure 2, the period near the cutoff time can be referred to as the present. We evaluate the model’s prediction accuracy across these time periods, denoted as $acc_{\text{past}}$, $acc_{\text{future}}$, and $acc_{\text{present}}$, respectively.
2.1.2 Define Temporal Bias and Temporal Degeneration
With the three time periods defined, we can define Temporal Bias and Temporal Degeneration for models.
Definition 2.2.
Temporal Bias In this paper, we define temporal bias as the preferences of LLMs over time.  [^1] Temporal bias are multifaceted in previous literature. For instance, can refer to the assumption of an incorrect sequence of events that misleads our reasoning about causality Dorsey (2021) or user preferences over time in point-of-interest recommendations Rahmani et al. (2022). Recently, biases have raised more challenges in pre-trained language models Yogarajan et al. (2023), particularly focusing on under-represented societies. Sarkar and Vafa (2024) examines lookahead bias in pre-trained language models, providing insights into how these biases affect model performance.
There exist two biases, namely Nostalgia Bias and Neophilia Bias, which reflect two contrasting tendencies in how these models might predict or generate text about the past, present, or future. A visual representation of these biases is provided in Fig. 1, where the trends associated with Nostalgia and Neophilia biases are plotted over time.
Definition 2.3.
Temporal Degeneration refers to the decline in a model’s performance in the future, that is, after it is released.
In conclusion, the illustration for the above concepts is in Figure 1.
@@ 2.2 Assessing Temporal Generalization
Based on definition in Sec. 2.1, we turn to assess temporal generalization through the lens of 1) I: compression intelligence in Sec. 2.2.1 and II): future event prediction in Sec. 2.2.2.
2.2.1 Case I: Compression Intelligence
@@
$\!BPC(T)\!=\!\frac{-\sum_{i=1}^{N}\log p(w_{i}|w_{1},\ldots,w_{i-1})}{\text{len-utf-8}(T)}$ (1)
@@
2.2.2 Case II: Future Event Prediction
LLMs could perform future event prediction, as seen in Bonde (2022) which employed GPT-3 in a few-shot setting to tackle binary questions from the Metaculus platform. Further, Zou et al. (2022) enriched the discourse by introducing a comprehensive dataset that includes forecasting questions alongside a news corpus. This approach enhances model training by incorporating varied and relevant data contexts. Moreover, Halawi et al. (2024) showcased how integrating news content can significantly improve the accuracy of predictions in real-time scenarios.
Building on this foundation, our research seeks to explore the temporal generalization capabilities of LLMs: how well they adapt and remain accurate as new data emerges and contexts evolve? As an example of using LLMs to predict political outcomes, we gather new factual data about future events, such as "What will President Biden’s approval rating be as of 7 June 2024, according to FiveThirtyEight?". To ensure objectivity in our predictions, answers are typically formatted as ranges within a prediction market scenario. For instance, President Biden’s approval rating might be categorized into intervals such as lower than 34%, at least 34.0%, but less than 36.0%, at least 36.0%, but less than 38.0%, up to 46.0% or higher.
@@ 2.2.2 Case II: Future Event Prediction
$Acc=\frac{N_{\text{correct}}}{N_{\text{total}}}$ (2)
@@ 3 Temporal Generalization through the Lens of Compression Intelligence
This section discusses how various models handle data from different periods, showcasing their performance stability or volatility.
Key Findings
  • In general, while the ranking of models varied slightly between datasets, the performance of individual models remained relatively stable over time within each dataset type, as can be seen in Figure 5 and 5. Here we made a discovery regarding the distinct characteristics of the Wikipedia and BBC datasets at 3.1.2.
  • Another interesting observation was made in the arXiv dataset shown in Figure  5, where a sharp decrease in BPC occurred around March 2024. This timing coincides with the peak in ChatGPT usage, as reported by Exploding Topics [^2] https://explodingtopics.com/blog/chatgpt-users, which saw over 1.8 billion visits in that month. In essence, it appears that humans may be adapting their writing to more closely resemble LLM outputs, rather than LLMs significantly enhancing their ability to model human language.
Figure 3: BPC on Wiki text
Figure 4: BPC on BBC news text
Figure 5: BPC on arXiv text
@@ 3.1.1 Protocol for Quantitative Measurement
Data sources Our system aggregates data from various platforms, including Yahoo for financial news, BBC for political insights, Reddit for discussions, Wikipedia for encyclopedic updates, arXiv for academic research, GitHub for software trends, and Quora for diverse questions and answers. Utilizing a Python-based crawling framework, it’s designed for adaptability and efficient data collection, with Playwright addressing navigation and specific site challenges. We also did some basic data pre-processing, due to space limitation, details in Appendix F. An overview of the data volume and the average entry length from a single crawl is provided in Appendix G. Experiment settings are listed in Appendix A.
@@ 3.1.2 Observation of Temporal Bias in Compression Intelligence
To comprehensively explore the temporal generalization capabilities of models, we plot the BPC curve over time.
In text data before release, a negative slope indicates the presence of Neophilia Bias. Conversely, a positive slope would suggest that the model’s performance is decreasing, which shows Nostalgia Bias.
ArXiv BPC has a huge decline after GPT is widely used, as mentioned in Key Findings. Models show stronger Nostalgia Bias when evaluated with BBC text than Wiki text. From the BPC charts, it’s not easy to spot differences between models. However, our calculations in Appendix I reveal variations in how well each model can adapt specific text over time.
Difference of Text Feature An insightful reviewer helped us found the reason of the different patterns of Wiki an BBC text. New Wiki texts with recent temporal markers do not necessarily imply that the content is new in a temporal context. For instance, in 2017-01, there was a Wiki article titled "History of India". However, the information within it was not new and could potentially have been sourced from other materials. Conversely, BBC texts are news articles and are highly time-sensitive.
To quantify the difference, we uniformly sampled approximately 500 data pieces from both the Wiki and BBC datasets (taking into account that news may also cover historical events) and asked gpt-4o to classify the content. The results are as follows:
  • Wiki: Historical: 97.83%, Others 333”Others” refers to content not associated with specific historical events, focusing instead on general, educational, and ever-relevant information. :2.17%.
  • BBC: Recent: 91.86%, Historical: 8.14%.
So, it is reasonable that Wiki text has a flat curve in BPC.
Claim the Value of Wiki Dataset Since the Wikipedia dataset is continuously updated, it is not easily subject to direct contamination. Although a substantial portion of Wiki pertains to old information, this content still represents valuable knowledge that we need. An ideal model should consistently perform well on such knowledge. (If a model shows a decline in performance on Wikipedia, it is likely that the model is overfitting.)
3.2 Temporal Degeneration
@@ 3.2.1 Protocol for Quantitative Measurement
Temporal Bias Index (TBI) For each model in a certain period, we record $BPC_{i}$ at $n$ time points, forming $n$ data points $(i,BPC_{i})$, where $i=1,\ldots,n$. Here the time interval between $BPC_{i}$ and $BPC_{i+1}$ is set as two month. We then fit a linear model:
$BPC_{i}=a*i+b$ (3)
then Temporal Bias Index (TBI) is defined as the coefficient $a$. TBI is crucial for understanding the direction and rate of change in model performance over time.
To quantify model temporal generalization, we computed the BPC for various models on the Wiki and BBC datasets and evaluated their performance using this index. Specifically, We established a base BPC using data from the six months prior to each model’s release. After that, we measured the changes in BPC at 3, 6, 9 and 12 months post-release relative to the base BPC; and assessed the overall trend in model performance using TBI. The specific results are detailed in Appendix H.
Table 1: Correlation between Model Attributes and Their TBI on Wiki and BBC text
@@ 3.2.2 Observation
Temporal generation pattern is mostly determined by the nature of data Models perform stably on Wikipedia but show more variation on BBC. This reveals their ability to handle well-structured and formal text like Wiki and insufficient adaptability to dynamic text like BBC text. The overall BPC for BBC texts shows an increasing trend, indicating gradual performance degeneration.
Individual models show considerable differences in performance on new data; for instance, Yi, and Baichuan2 series adapt relatively well to new texts, while Phi-1.5 and OPT series exhibit poorer generalization (see Appendix I). These findings indicate that in dynamic environments, different models have significantly varying adaptation capabilities.
Model performance positively correlates with performance decay. (see Tab. 1). There is a positive correlation between model attributes (such as model size and release date) and performance decay. A positive BPC slope indicates an increase in BPC over time, reflecting performance degeneration.
In the Wiki dataset, larger models deteriorate faster (correlation coefficient of 0.504), suggesting a trade-off between capability and flexibility. In more dynamic environments (such as the BBC dataset), the impact of model size on performance decay is smaller. Additionally, newer models show higher correlations, meaning that despite their advanced capabilities, they may deteriorate faster. These results emphasize the importance of designing models that can maintain robustness over time and across various data environments. As machine learning applications grow in complexity and scale, temporal generalization will become increasingly critical.
4 Temporal Generalization through the Lens of Future Event Prediction
Advantages of Future Event Prediction over Compression Intelligence Compression Intelligence is limited by its inapplicability to closed-source models, where key indicators like logits or losses are inaccessible. This exclusion narrows comparative analysis and hinders a comprehensive understanding of the language technology landscape. Additionally, language likelihood can be biased by models finely tuned to specific formats or domains, skewing overall evaluation and presenting a distorted view of a model’s general performance.
This motivated us to research temporal generalization and bias from the perspective of future prediction.
In this section, we investigate temporal generalization (in Sec. 4.1) and biases (Sec. 4.2) through the Lens of Future Event Prediction.
Key Findings
  • Models prefer older knowledge, with non-uniform knowledge distribution skewed towards the past. This poses a risk as users expect up-to-date knowledge close to the model’s release date.
  • Within model series such as Claude,Gemini, and GPT, newer versions are becoming more powerful but also experience more rapid performance decline. Miniature models significantly underperform compared to their full -sized counterparts in temporal tasks.
  • Newer models generally exhibit stronger performance. Closed-source models tend to perform slightly better at the beginning, yet open-source models demonstrate superior generalization capabilities.